# Lessons Learned So Far From Verifying the Rust Standard Library (work-in-progress)

### v2

@@ Abstract.
Although Rust primarily intends to be a safe programming language that excludes undefined behaviour, it provides its users with the escape hatch of unsafe Rust, allowing them to circumvent some of its strong compile-time checks. This additional freedom has some advantages, including potentially more efficient code, which is one of the main reasons why unsafe code is used extensively throughout Rust’s standard library. However, because unsafe code also re-opens the door to undefined behaviour, Amazon has convened a community to verify the safety of the standard library, and in particular the unsafe code contained therein. Given that this effort is done in public and open-sourced, we have access to a wealth of information on how people are verifying the standard library, as well as what is currently possible and what still appears to be beyond the state of the art for verified software.
@@ 1. Introduction
For these reasons, Amazon has proposed a community-driven effort to verify Rust’s standard library (Kumar et al., 2024), which it targets in a segmented way, via challenges [^2] https://model-checking.github.io/verify-rust-std/intro.html. These challenges focus primarily on safety properties rather than functional correctness. One reason for the focus on safety is that, while there is a well-understood set of safety properties to check (given that we have an exhaustive list of possible sources of undefined behaviour [^3] https://doc.rust-lang.org/reference/behavior-considered-undefined.html), there is no real consensus on the scope for functional (or domain-specific) correctness. We also have earlier work that commented specifically on this Amazon-led effort, including (Blanc and Lam, 2024), which discusses what it means to verify the standard library, along with some suggestions on how best to do so.
In this paper, we discuss what we have learned so far from our own contributions to this effort (namely, verifying transmute and its uses). Alongside this, we also present what we have learned from studying how the other challenges are being solved.
@@ 2.1. Verification progress
As shown in Table 1, to date, 27 challenges have been posted, 9 of which have accepted solutions. The remaining 18 challenges are still open, and 3 of these have seen nontrivial engagement. The challenges target parts of the standard library’s three main crates (core, alloc, and std) that might be particularly vulnerable to memory safety issues (such as callers of transmute and raw pointer arithmetic operations) as well as data races (such as concurrency primitives). The rewards for completing the challenges, currently ranging from 5000 to 25000 USD, are based on the perceived difficulty of completion, as determined by the committee [^5] This committee consists primarily of Amazon staff, along with a few other invited members from academia and industry. They are in charge of reviewing newly proposed challenges, challenge solutions, and new tools. for this verification effort.
@@ 2.2. The tools used
The first tool approved for use in this effort was Kani, and to date, 6 out of 9 of the completed challenges use Kani. Since then, three other tools have been approved by the committee and integrated into the effort’s Continuous Integration workflow, with three more in the process of being approved. We present them in Table 2, and briefly discuss the properties of these tools and the potential for other tools.
@@ Abstract
Kani (VanHattum et al., 2022) uses bounded model checking as implemented in CBMC (Clarke et al., 2004). Kani also has experimental support for loop invariants, which enable it to go beyond some of the limitations of bounded proof.
To give an example of Kani’s use, in Listing 2, we have an integer doubling function that returns None if the input is greater than half the maximum i32, so as to avoid a potential integer overflow (note: this is just the specification of the hypothetical function, overflowing is not itself undefined behaviour). To check that double does not overflow and indeed returns None when its input is greater than half the maximum i32, we can write a ‘proof harness’, i.e., a separate function analogous to a test harness, but for proving that a property holds, rather than for testing. See Listing 3 for an example.
@@ 6 Some(x * 2)
A second backend for Kani is the goto-transcoder, which generates inputs suitable for the ESBMC tool (Cordeiro et al., 2009) (rather than CBMC, as usually targetted by Kani). Using the goto-transcoder adds support for $k$-induction and SMT solvers. In principle, this could result in more powerful verification (more cases verified) using less CPU time and memory, but, for now, no challenges have been solved using goto-transcoder rather than Kani.
@@
VeriFast (Jacobs et al., 2015) represents memory using separation logic and symbolically executes methods to verify them. In this context, VeriFast users effectively create annotated copies of standard library code augmented with VeriFast-specific annotations. Such annotations notably include loop invariants and inductive predicates, which VeriFast uses to perform unbounded verification over loops and dynamically-sized types, respectively (see Section 5 for further discussion on boundedness). These annotated functions are symbolically executed, and the resulting verification condition is passed to an SMT solver. VeriFast’s use of separation logic allows for the expression of richer properties than other tools (like Kani), particularly ones complicated by the problem of aliasing (e.g., linked-list properties). There is a provision for automatically bringing the VeriFast changes up to date with changes to the standard library. At the time of writing, VeriFast is the only tool besides Kani (and the randomized testing for Challenge 15) to have been used in an accepted challenge solution (2 of the 9 accepted solutions).
@@
Another approved tool is Flux (Lehmann et al., 2023), a refinement type checker for Rust. At the time of writing, Flux has not yet solved any challenges. Compared to VeriFast, it aims to be more lightweight, which it does by restricting predicates to first-order logic only (as opposed to VeriFast’s separation logic), which does come with the disadvantage of making some properties harder to specify. Unlike Kani and goto-transcoder, it performs unbounded verification (it automatically infers loop invariants, and it allows users to specify invariants for even dynamically-sized types). However, it currently offers only limited support for unsafe code (e.g., it cannot track values written through pointers).
@@
There are also three tools in the process of approval, namely KMIR, Creusot (Denis et al., 2022), and RAPx [^7] https://github.com/Artisan-Lab/RAPx. KMIR is developed by the company Runtime Verification, and uses the K framework [^8] https://kframework.org to define operational semantics for Rust’s Middle Intermediate Representation. KMIR performs symbolic execution on the IR and uses reachability logic to verify needed conditions; it claims little dependence on SAT or SMT solvers. Creusot is similar to VeriFast in that it is a deductive verifier, but it only uses first-order logic as opposed to separation logic, which again means that it can be fully automated and lightweight, but in exchange, it loses out on expressiveness. It still achieves unboundedness, via loop invariants and inductive properties. Finally, RAPx is a tool that performs contract-based abstract interpretation, and can automatically infer safety conditions for unsafe APIs. As is typical of similar static tools, it over-approximates, but achieves soundness in return.
@@ 2.3. Suggested tools for remaining challenges
In Table 3, we present the main verification obstacles for each remaining challenge, based on the above criteria. We further list which of the accepted tools are capable of completing the challenges [^9] This is purely based on the identified verification obstacles. In reality, it is possible that some of the tools listed cannot solve the challenges for reasons that we are unaware of.. We further note that while hybrid solutions (i.e., ones involving a combination of tools) have not yet been used or discussed, there certainly would be some value in using them (e.g., for Challenge 16, we could use VeriFast specifically for any function that involves unbounded or generic-typed reasoning, and then the more lightweight Kani for the others).
@@ 5. Discussion
This verification effort is challenge-driven. Amazon staff have proposed almost all of the challenges so far, but our experience is that it is possible to propose challenges from outside as well. At least 2 of the existing challenges did not originate from Amazon. As summarized in Section 2.2, non-Amazon groups appear to be more keen to contribute tools (6/7) than challenges.

### v3

# Lessons Learned So Far From a Community Effort to Verify the Rust Standard Library (work-in-progress)

@@ Abstract.
Although Rust primarily intends to be a safe programming language that excludes undefined behaviour, it provides its users with the escape hatch of unsafe Rust, allowing them to circumvent some of its strong compile-time checks. This additional freedom has some advantages, including potentially more efficient code, which is one of the main reasons why unsafe code is used extensively throughout Rust’s standard library. However, because unsafe code also re-opens the door to undefined behaviour, Amazon has convened a community to verify the safety of the standard library, and in particular the unsafe code contained therein. Given that this effort is done in public, is open-sourced, and has seen significant participation (from at least 50 different contributors), we have access to a wealth of information on how people are verifying the standard library, as well as what is currently possible and what still appears to be beyond the state of the art for verified software.
@@ 1. Introduction
For these reasons, Amazon has proposed a community-driven endeavour to verify Rust’s standard library (Kumar et al., 2024). As part of ensuring broad community support, Amazon has engaged the Rust Foundation, which has agreed to host the endeavour. The endeavour aims to verify the standard library in a segmented way, via challenges [^2] https://model-checking.github.io/verify-rust-std/intro.html. These challenges focus primarily on safety properties rather than functional correctness. One reason for the focus on safety is that, while there is a well-understood set of safety properties to check (given that we have an exhaustive list of possible sources of undefined behaviour [^3] https://doc.rust-lang.org/reference/behavior-considered-undefined.html), there is no real consensus on the scope for functional (or domain-specific) correctness. We also have earlier work that commented specifically on this Amazon-led effort, including (Blanc and Lam, 2024), which discusses what it means to verify the standard library, along with some suggestions on how best to do so.
In this paper, we discuss what we have learned so far from our own contributions to this effort (namely, verifying transmute and its uses). Alongside this, we also present what we have learned from studying how the other challenges are being solved. It should be noted that we represent only a small fraction of this larger community, which includes over 50 contributors. Indeed, the purpose of this paper is twofold: (1) to introduce this verification effort to an even broader audience, inciting additional participation; and, (2) to provide novel observations to help people newly interested in this effort, as well as veteran contributors.
@@
Disclaimer.
As previously stated, we represent only a small portion of the entire community working on this Amazon-led verification effort. In writing this paper, we explicitly do not claim credit for the vast majority of the effort—we only claim credit for having contributed a solution to 1 of the 27 challenges and proposing 2 of the challenges. Rather, we are putting forward some observations we have made from our vantage point from the periphery of the community. Our goals are to increase awareness of and interest in the effort. We thank all contributors to this effort for their significant efforts so far, and especially Amazon and its staff for their leadership role.
@@ 2.1. Verification progress
As shown in Table 1, to date, 27 challenges have been posted, 9 of which have accepted solutions. The remaining 18 challenges are still open, and 3 of these have seen non-trivial engagement. The challenges target parts of the standard library’s three main crates (core, alloc, and std) that might be particularly vulnerable to memory safety issues (such as callers of transmute and raw pointer arithmetic operations) as well as data races (such as concurrency primitives). Almost all challenges (24/27) have been put forth by Amazon, whereas 3 have been suggested by external community members (Challenge 15 by people at Cryspen, and Challenges 26 and 27 by us from the University of Waterloo). The rewards for completing the challenges, currently ranging from 5000 to 25000 USD, are based on the perceived difficulty of completion, as determined by the committee [^5] This committee consists primarily of Amazon staff, along with a few other invited members from academia and industry. They are in charge of reviewing newly proposed challenges, challenge solutions, and new tools. for this verification effort. Moreover, challenge solutions are peer-reviewed, almost entirely by Amazon staff [^6] The most prolific reviewers in terms of pull requests reviewed are Michael Tautschnig (231), Carolyn Zech (132), Felipe R. Monteiro (88), Celina G. Val (60), Zyad Hassan (40), Qinheping Hu (29), Thanh Nguyen (20), Rahul Kumar (14), and Jaisurya Nanduri (13), all from AWS.. The efforts of Amazon staff have been instrumental in the success of this endeavour to date.
@@ 2.2. The tools used
The first tool approved for use in this effort was Kani (VanHattum et al., 2022), and to date, 6 out of 9 of the completed challenges use Kani. Since then, three other tools have been approved by the committee and integrated into the effort’s Continuous Integration workflow, with three more in the process of being approved. We present them in Table 2, and briefly discuss the properties of these tools and the potential for other tools.
@@ Abstract
Kani (VanHattum et al., 2022), by Amazon, uses bounded model checking as implemented in CBMC (Clarke et al., 2004). Kani also has experimental support for loop invariants, which enable it to go beyond some of the limitations of bounded proof.
To give an example of Kani’s use, in Listing 2, we have an integer doubling function that returns None if the input is greater than half the maximum i32, so as to avoid a potential integer overflow (note: this is just the specification of the hypothetical function; overflowing is not itself undefined behaviour). To check that double does not overflow and indeed returns None when its input is greater than half the maximum i32, we can write a ‘proof harness’, i.e., a separate function analogous to a test harness, but for proving that a property holds, rather than for testing. See Listing 3 for an example.
@@ 6 Some(x * 2)
A second backend for Kani is the goto-transcoder, by Rafael Sá Menezes, which generates inputs suitable for the ESBMC tool (Cordeiro et al., 2009) (rather than CBMC, as usually targetted by Kani). Using the goto-transcoder adds support for $k$-induction and SMT solvers. In principle, this could result in more powerful verification (more cases verified) using less CPU time and memory, but, for now, no challenges have been solved using goto-transcoder rather than Kani.
@@
VeriFast (Jacobs et al., 2015), by Jacobs et al., represents memory using separation logic and symbolically executes methods to verify them. In this context, VeriFast users effectively create annotated copies of standard library code augmented with VeriFast-specific annotations. Such annotations notably include loop invariants and inductive predicates, which VeriFast uses to perform unbounded verification over loops and dynamically-sized types, respectively (see Section 5 for further discussion on boundedness). These annotated functions are symbolically executed, and the resulting verification condition is passed to an SMT solver. VeriFast’s use of separation logic allows for the expression of richer properties than other tools (e.g. Kani), particularly ones complicated by the problem of aliasing (e.g., linked-list properties). There is a provision for automatically bringing the VeriFast changes up to date with changes to the standard library. At the time of writing, VeriFast is the only tool besides Kani (and the randomized testing for Challenge 15) to have been used in an accepted challenge solution (2 of the 9 accepted solutions).
@@
Another approved tool is Flux (Lehmann et al., 2023) (by Lehmann et al.), a refinement type checker for Rust. At the time of writing, Flux has not yet solved any challenges. Compared to VeriFast, it aims to be more lightweight, which it does by restricting predicates to first-order logic only (as opposed to VeriFast’s separation logic). This restriction does come with the disadvantage of making some properties harder to specify. Unlike Kani and goto-transcoder, it performs unbounded verification (it automatically infers loop invariants, and it allows users to specify invariants for even dynamically-sized types). However, it currently offers only limited support for unsafe code (e.g., it cannot track values written through pointers).
@@
There are also three tools in the process of approval, namely KMIR, Creusot (Denis et al., 2022), and RAPx [^8] https://github.com/Artisan-Lab/RAPx. KMIR is developed by the company Runtime Verification, and uses the K framework [^9] https://kframework.org to define operational semantics for Rust’s Middle Intermediate Representation. KMIR performs symbolic execution on the IR and uses reachability logic to verify needed conditions; it claims little dependence on SAT or SMT solvers. Creusot, by Denis et al., is similar to VeriFast in that it is a deductive verifier, but it only uses first-order logic as opposed to separation logic. This means that it can be fully automated and lightweight, but in exchange, it loses out on expressiveness. It still achieves unboundedness, via loop invariants and inductive properties. Finally, RAPx (by researchers at Fudan University) is a tool that performs contract-based abstract interpretation, and can automatically infer safety conditions for unsafe APIs. As is typical of similar static tools, it over-approximates, but achieves soundness in return.
@@ 2.3. Suggested tools for remaining challenges
In Table 3, we present the main verification obstacles for each remaining challenge, based on the above criteria. We further list which of the accepted tools are capable of completing the challenges [^10] This is purely based on the identified verification obstacles. In reality, it is possible that some of the tools listed cannot solve the challenges for reasons that we are unaware of.. We further note that while hybrid solutions (i.e., ones involving a combination of tools) have not yet been used or discussed, there certainly would be some value in using them (e.g., for Challenge 16, one could use VeriFast specifically for any function that involves unbounded or generic-typed reasoning, and then the more lightweight Kani for the others).
@@ 5. Discussion
This verification effort is challenge-driven. Amazon staff have proposed almost all of the challenges so far, but our experience is that it is possible to propose challenges from outside as well. As stated eariler, 3 of the existing challenges did not originate from Amazon. As summarized in Section 2.2, non-Amazon groups appear to be more keen to contribute tools (6/7) than challenges.
@@ 6. Conclusion
We encourage researchers and practitioners interested in formal methods to join the standard library verification effort initiated by Amazon, and we hope that this paper also serves as a useful starting point for anyone wishing to do so.